OpenStudy (anonymous):

In the 2004 Summer Olympics Games in Athens, Greece, the United States athletes won 2 more than 3 times the number of gold medals won by French athletes. The United States won 24 more gold medals than the French. Write and solve an equation to determine the number of medals won by the French

OpenStudy (anonymous):

U = 2+3f U = 24 + f U represents the medals won by the United States, and f represents those won by France. From there, plug in the equations using substitution and get 2 + 3f = 24 + f And so, f = 11.